Night report 2017-05-18

Night report 2017-05-18

Observers: Alessandro Razza, Felipe Olivares.

TAT on duty: Michel Dennefeld, Cosimo Inserra, Giorgos Leloudas.

Support team: Lluis Galbany.

Daytime calibrations: PESSTO_calib, PESSTO_bias

Owing to bad weather conditions skyflats were not obtained.

Weather conditions: 

Snow-frozen domes, Humidity > 90%.

Dome closed the whole night for the weather conditions.